This is my question--From an urn containing 7 red and 4 green balls, in how many ways can we draw a set of 5 balls of which 3 are red and 2 are green?
Answer by stanbon(75887)

From an urn containing 7 red and 4 green balls, in how many ways can we draw a set of 5 balls of which 3 are red and 2 are green?
---------------
# of ways to draw 3 red of 7 is 7C3= [7*6*5]/[1*2*3]= 35
# of ways to draw 2 green of 4 is 4C2=[4*3}/[1*2]=6
----------
# of ways to do both is 35*6 = 210 ways
===========
Cheers,
Stan H.